# U-MARVEL: Unveiling Key Factors for Universal Multimodal Retrieval via Embedding Learning with MLLMs

### v1

@@ Abstract
Universal multimodal retrieval (UMR), which aims to address complex retrieval tasks where both queries and candidates span diverse modalities, has been significantly advanced by the emergence of MLLMs. While state-of-the-art MLLM-based methods in the literature predominantly adopt contrastive learning principles, they often differ in their specific training recipes. Despite their success, the mechanisms underlying their retrieval capabilities remain largely unexplored, potentially resulting in suboptimal performance and limited generalization ability. To address these issues, we present a comprehensive study aimed at uncovering the key factors that drive effective embedding learning for UMR using MLLMs. We begin by implementing a general MLLM-based embedding learning pipeline, and systematically analyze the primary contributors to high-performing universal retrieval systems. Based on this, we explore various aspects of the details in embedding generation and training strategies, including progressive transition, hard negative mining and re-ranker distillation. Notably, our findings reveal that often-overlooked factors can have a substantial impact on model performance. Building on these discoveries, we introduce a unified framework termed U-MARVEL (Universal MultimodAl RetrieVal via Embedding Learning), which outperforms state-of-the-art competitors on the M-BEIR benchmark by a large margin in supervised settings, and also exihibits strong zero-shot performance on several tasks such as composed image retrieval and text-to-video retrieval. These results underscore the generalization potential of our framework across various embedding-based retrieval tasks. Code is available at https://github.com/chaxjli/U-MARVEL
@@ 1 Introduction
Multimodal information retrieval [37] has emerged as a pivotal research direction at the intersection of multimodal understanding and information retrieval. It serves as a cornerstone for a wide range of modern AI applications, including retrieval-augmented generation (RAG) [23, 10], text-image retrieval [3, 60], and visual question answering (VQA) [8, 9]. While existing methods such as CLIP [42], BLIP [28, 27], and CoCa [56] have demonstrated impressive performance in cross-modal retrieval, they often struggle to address the diverse and complex requirements encountered in real-world scenarios.
To tackle these challenges, the research community has increasingly focused on universal multimodal retrieval (UMR) [51], depicted in Fig. 1(a), which aims to develop unified, instruction-guided retrievers capable of handling diverse retrieval tasks across multiple modalities. Recent advancements in this area, such as LamRA [36], MM-Embed [30], GME [62], UniME [15] and VLM2VEC [24], have shown promising results on UMR powered by MLLMs. However, most of the existing approaches often directly adapt MLLMs to embedding tasks without systematically examining training schemes tailored for MLLM-based embedding models. Consequently, the impact of design decisions on retrieval performance remains inadequately understood, highlighting the need for a more systematic investigation—a gap that motivates our study.
@@
Motivated by these insights, we present U-MARVEL, a unified framework for learning universal multimodal embeddings tailored to UMR tasks. Our method achieves significant improvements over state-of-the-art UMR approaches on the M-BEIR benchmark [51] in supervised settings, and also demonstrates strong zero-shot generalization abilities on tasks such as text-to-video retrieval and composed image retrieval.
@@ 2 Related Work
Large language models (LLMs) [1, 54, 32, 45], including multimodal LLMs [61, 55], have achieved remarkable success across a wide range of general-purpose tasks. For instance, LLaVA [34] integrates image and text processing for tasks like visual question answering, image description, and visual reasoning. The Qwen-VL series [47, 2] enhances this by supporting flexible image resolutions and extending capabilities to video understanding through temporal modeling and localization. InternVL3 [63] unifies multimodal and linguistic learning by leveraging both multimodal data and text corpora in a single pre-training stage.
@@
Multimodal retrieval aims to align diverse modalities (e.g., text, images, or their combinations) within a shared embedding space for semantic similarity matching. Early works [42, 28, 56] leveraged large-scale contrastive learning to align image-text pairs, enabling zero-shot classification and cross-modal retrieval, but struggled with complex multimodal inputs. Recent advancements have addressed these limitations through innovative approaches based on MLLMs [22, 51, 24, 25]. LamRA [36] adapts MLLMs into retrievers using lightweight LoRA modules with joint pointwise and listwise reranking; MM-Embed [30] improves alignment via MLLM fine-tuning and modality-aware hard negative mining; GME [62] develops a unified multimodal retrieval framework leveraging MLLMs with synthetic data augmentation; and UniME [15] combines textual knowledge distillation from LLM teachers with hard negative-enhanced instruction tuning, achieving robust cross-modal alignment for diverse retrieval tasks. While these works provide a strong foundation for UMR, the design space for MLLM-based embedding models remains underexplored. Unlike LLMs that generate outputs via autoregressive decoding, embedding models encode features of the entire input sequence, necessitating specialized strategies for feature extraction and tailored training methodologies. This work addresses these gaps through a systematic investigation of the design decisions for MLLM-based embedding models.
@@ 3 Recipe for Building U-MARVEL
In this section, we conduct a comprehensive study to identify the key factors that drive effective embedding learning for UMR through MLLMs. Following the mainstream approaches, we first establish a simple pipeline. Unless stated otherwise, we finetune the pre-trained Qwen2-VL-7B-Instruct [47] as the embedding model with LoRA [19] by contrastive learning on M-BEIR dataset [51]. Details are given in Appendix 3. We then explore three major axes of design decisions in the following subsections:
@@ 3.1.1 Embedding Extraction
We first investigate the method for extracting embeddings of the token sequence of multimodal queries using MLLMs. Following prior approaches, the instruction is directly concatenated with the query. Existing methods can basically be divided into two categories based on the way embeddings are extracted. (1) Last token: previous works [36, 24, 62, 22, 15] typically employ a prompt for compression. Generally, for multimodal input, the prompt is "<image><text>…<image><text> Summarize above image and sentence in one word: <emb>", where <image> and <text> are the placeholders. In this way, the feature of last token (i.e., <emb>) is used as the output embedding of the input sequence. (2) Mean token: an alternative method, adopted by MM-Embed [30], transforms the unidirectional attention mechanism into a bidirectional one and applies mean pooling to the features from the last hidden states to derive the embedding for the entire sequence.
@@
| Bidirectional | Mean token | ("in one word") | Avg. | Avg. | |
| 0 | Causal | Last token | ✓ | 56.6 | 54.8 |
| 1 | Bidirectional | Last token | ✓ | 55.5 | 52.6 |
| 2 | Causal | Mean token | ✓ | 33.7 | 27.6 |
| 3 | Bidirectional | Mean token | ✓ | 51.7 | 46.1 |
| 4 | Bidirectional | Mean token | ✕ | 57.2 | 55.2 |
@@
In conclusion, generating embeddings using bidirectional attention combined with mean pooling from the sequence demonstrates superior performance compared to the commonly employed approach of combining compression prompts with the last token mechanism. This difference may stem from the last token embedding being affected by recency bias, leading to an excessive reliance on the output of the final token. Notably, this finding aligns with the conclusions of NV-Embed [26], yet diverges from those presented in GME [62], offering a distinct perspective on architectural design and performance optimization.
@@ 3.1.2 Instruction Integration
Motivated by prior studies on text embeddings [26, 4], we mask out the instruction tokens during the mean pooling process, as these tokens have already influenced the output features through self-attention mechanisms. Experimental results show that ID-0 achieves a 0.1%/0.3% improvement compared to ID-1 and 2.0%/12.1% improvement compared to ID-2 in Table 3. We hypothesize that this improvement arises primarily from that: the query inherently incorporates instruction information during forward propagation thought bidirectional self-attention mechanism. Therefore, by filtering out instruction-related information and focusing on feature comparisons between the query and candidate, the approach mitigates calculation bias.
@@ 3.1.3 Progressive Transition
We set a baseline by vanilla instruction tuning on multimodal retrieval data from the M-BEIR dataset, presented in Table 3 under ID-0. Following prevalent approaches [36, 22], we use text-only retrieval data for pre-training, which improves performance on various retrieval tasks. Training on the NLI dataset [13] significantly enhances our model’s performance, as indicated by ID-1 in Table 3. Despite these improvements, a notable gap persists when transitioning from single-modality, single-instruction tasks to multimodal, multi-instruction tasks.
To address this gap, we propose a progressive approach incorporating an additional pre-training step using text-image retrieval data from the CC3M dataset [43]. This step further enhances performance, as demonstrated by ID-2 in Table 3. Since the LLM was originally trained with causal attention, switching to bidirectional attention degrades text-visual encoder alignment. We speculate that text-based fine-tuning strengthens the text encoder’s semantic representation, while image-text pairs further align the text and visual encoders. During experiments, we train the model with NLI using unidirectional InfoNCE loss, followed by CC3M using bidirectional InfoNCE loss, pairing each query with its correct match (positive sample) while other items in the batch serve as negative samples.
Therefore, we formally propose the progressive transition method, adopting a stepwise training strategy to adapt decoder-only MLLMs for embedding generation tasks. Following the principle of advancing from simpler to more complex tasks, this phased approach ensures a smooth transition through three key steps: (1) Text Retrieval Adaptation: Learning retrieval tasks using text-only datasets to establish foundational capabilities. (2) Cross-modal Alignment: Advancing to cross-modal retrieval tasks through text-image paired data. (3) Instruction Tuning for Multimodal Retrieval: Mastering instruction-guided multimodal retrieval tasks with comprehensive training on multimodal datasets.
@@ 3.2.1 Interactions Among Large Batch Size, Learning Rate and Temperature
Our experimental results dispute the widely held belief that "increasing batch size directly improves performance", indicating that this assumption requires closer examination. By analyzing the results presented in Table 4, we identify three key findings as follows. (1) Experiments with ID-0, ID-2, and ID-4 indicate that increasing the batch size does improve model performance; however, the performance gains tend to plateau as the batch size continues to grow, as observed in comparisons between ID-5 and ID-6. (2) Consistent with prior research [14] that highlights the necessity of scaling the learning rate (lr) alongside batch size, our contrastive learning experiments reveal similar trends. Specifically, simply increasing the batch size without adjusting the learning rate results in marginal improvements (ID-0 vs. ID-1). Optimal learning rate scaling approximately follows the square root of the batch size increase ratio, as demonstrated in comparisons between ID-0 and ID-2. (3) Our investigation into the effects of temperature parameters reveals that employing a learnable temperature significantly enhances performance. This is evident from comparing ID-4 with ID-7 and ID-6 with ID-8. While existing methods such as LamRA [36] and MM-Embed [30] utilize a fixed temperature parameter, we argue that the learnable temperature parameter in contrastive learning dynamically adjusts the sharpness of the probability distribution, enabling the model to optimize this critical hyperparameter during training. By adapting to different training stages, dataset characteristics, and representation quality, it simplifies hyperparameter tuning, improves gradient properties, balances competing objectives, and enhances overall performance. This innovation makes contrastive learning more robust and effective across diverse applications.
It is important to note that all experiments were conducted within a single training epoch, meaning that larger batch sizes corresponded to fewer training steps. Furthermore, increasing the number of epochs provided negligible additional benefits when hyperparameters were appropriately configured.
@@ 3.2.2 Continual Training with Hard Negative Mining
The initial random negative sampling yields basic discriminative ability but struggles with challenging false positives: negatives that exhibit high semantic similarity (measured by similarity scores) yet are incorrect matches. Recent studies have employed hard negative mining [30, 25] or hard negative reweighting in loss computation [17, 41, 64] to address this issue. In this work, we explore related settings. However, as shown in Table 5 for ID-0, directly selecting the top- hard negatives proved to be excessively challenging, resulting in complete failure. Additionally, incorporating in-batch negatives results in a performance decline compared to the baseline in this table, as observed by ID-1. We hypothesize that some hard negatives are false negatives, which can adversely affect model training. To address this issue, we propose a straightforward filtering method that eliminates hard negatives with scores exceeding a predefined threshold. This approach led to a significant improvement in model performance, as demonstrated by ID-2.
We apply a hard negative mining strategy as follows: (1) Feature Extraction: Using the model from the previous training stage, we extract features for all queries and candidates, compute similarity scores (excluding positive matches), and rank candidates to identify the most challenging negatives per query; (2) Filtered Hard Negative: For each query, we filter out negative samples that exceed a predefined threshold (considered as false negatives) and then select the top- negative samples as hard negatives; (3) Balanced Training: To prevent model convergence difficulties from excessive hard negatives and to accelerate training, we select a suitable -value and mix these hard negatives with other in-batch negatives. We then perform continual finetuning by InfoNCE.

### v2

@@ Abstract
Universal multimodal retrieval (UMR), which aims to address complex retrieval tasks where both queries and candidates span diverse modalities, has been significantly advanced by the emergence of MLLMs. While state-of-the-art MLLM-based methods in the literature predominantly adopt contrastive learning principles, they often differ in their specific training recipes. Despite their success, the mechanisms underlying their retrieval capabilities remain largely unexplored, potentially resulting in suboptimal performance and limited generalization ability. To address these issues, we present a comprehensive study aimed at uncovering the key factors that drive effective embedding learning for UMR using MLLMs. We begin by implementing a general MLLM-based embedding learning pipeline, and systematically analyze the primary contributors to high-performing universal retrieval systems. Based on this, we explore various aspects of the details in embedding generation and training strategies, including progressive transition, hard negative mining and re-ranker distillation. Notably, our findings reveal that often-overlooked factors can have a substantial impact on model performance. Building on these discoveries, we introduce a unified framework termed U-MARVEL (Universal MultimodAl RetrieVal via Embedding Learning), which outperforms state-of-the-art competitors on the M-BEIR benchmark by a large margin in supervised settings, and also exhibits strong zero-shot performance on several tasks such as composed image retrieval and text-to-video retrieval. These results underscore the generalization potential of our framework across various embedding-based retrieval tasks. Code is available at https://github.com/chaxjli/U-MARVEL.
@@ 1 Introduction
Multimodal information retrieval (Liu et al., 2022) has emerged as a pivotal research direction at the intersection of multimodal understanding and information retrieval. It serves as a cornerstone for a wide range of modern AI applications, including retrieval-augmented generation (RAG) (Jiang et al., 2023; Cong et al., 2023), text-image retrieval (Baldrati et al., 2023; Zhang et al., 2024b), and visual question answering (VQA) (Chow et al., 2024; Chun et al., 2021). While existing methods such as CLIP (Radford et al., 2021), BLIP (Li et al., 2022; 2023a), and CoCa (Yu et al., 2022) have demonstrated impressive performance in cross-modal retrieval, they often struggle to address the diverse and complex requirements encountered in real-world scenarios, ranging from fine-grained instruction following to multi-turn interleaved interactions.
To tackle these challenges, the research community has increasingly focused on universal multimodal retrieval (UMR) (Wei et al., 2024), depicted in Fig. 1(a), which aims to develop unified, instruction-guided retrievers capable of handling diverse retrieval tasks across multiple modalities. Recent advancements in this area, such as LamRA (Liu et al., 2024b), MM-Embed (Lin et al., 2024), GME (Zhang et al., 2024d), UniME (Gu et al., 2025) and VLM2VEC (Jiang et al., 2024b), have shown promising results on UMR powered by MLLMs. However, most of the existing approaches often directly adapt MLLMs to embedding tasks without systematically examining training schemes tailored for MLLM-based embedding models. Consequently, the impact of design decisions on retrieval performance remains inadequately understood, highlighting the need for a more systematic investigation—a gap that motivates our study.
@@
Motivated by these insights, we present U-MARVEL, a unified framework for learning universal multimodal embeddings tailored to UMR tasks. Our method achieves significant improvements over state-of-the-art UMR approaches on the M-BEIR benchmark (Wei et al., 2024) in supervised settings, and also demonstrates strong zero-shot generalization abilities on tasks such as text-to-video retrieval and composed image retrieval.
@@ 2 Related Work
Large language models (LLMs) (Achiam et al., 2023; Yang et al., 2024; Liu et al., 2024a; Touvron et al., 2023), including multimodal LLMs (Zhang et al., 2024c; Yao et al., 2024), have achieved remarkable success across a wide range of general-purpose tasks. For instance, LLaVA (Liu et al., 2023a) integrates image and text processing for tasks like visual question answering, image description, and visual reasoning. The Qwen-VL series (Wang et al., 2024a; Bai et al., 2025) enhances this by supporting flexible image resolutions and extending capabilities to video understanding through temporal modeling and localization. InternVL3 (Zhu et al., 2025) unifies multimodal and linguistic learning by leveraging both multimodal data and text corpora in a single pre-training stage.
@@
Multimodal retrieval aims to align diverse modalities (e.g., text, images, or their combinations) within a shared embedding space for semantic similarity matching. Early works (Radford et al., 2021; Li et al., 2022; Yu et al., 2022) leveraged large-scale contrastive learning to align image-text pairs, enabling zero-shot classification and cross-modal retrieval, but struggled with complex multimodal inputs. Recent advancements have addressed these limitations through innovative approaches based on MLLMs (Jiang et al., 2024a; Wei et al., 2024; Jiang et al., 2024b; Lan et al., 2025). LamRA (Liu et al., 2024b) adapts MLLMs into retrievers using lightweight LoRA modules with joint pointwise and listwise reranking; MM-Embed (Lin et al., 2024) improves alignment via MLLM fine-tuning and modality-aware hard negative mining; GME (Zhang et al., 2024d) develops a unified multimodal retrieval framework leveraging MLLMs with synthetic data augmentation; and UniME (Gu et al., 2025) combines textual knowledge distillation from LLM teachers with hard negative-enhanced instruction tuning, achieving robust cross-modal alignment for diverse retrieval tasks. While these works provide a strong foundation for UMR, the design space for MLLM-based embedding models remains underexplored. Unlike LLMs that generate outputs via autoregressive decoding, embedding models encode features of the entire input sequence, necessitating specialized strategies for feature extraction and tailored training methodologies. This work addresses these gaps through a systematic investigation of the design decisions for MLLM-based embedding models.
@@ 3 Recipe for Building U-MARVEL
In this section, we conduct a comprehensive study to identify the key factors that drive effective embedding learning for UMR through MLLMs. Following the mainstream approaches, we first establish a simple pipeline. Unless stated otherwise, we finetune the pre-trained Qwen2-VL-7B-Instruct (Wang et al., 2024a) as the embedding model with LoRA (Hu et al., 2022) by contrastive learning on M-BEIR dataset (Wei et al., 2024). Specifically, the InfoNCE loss (Oord et al., 2018) is selected as our training objective:
where , , represent the embeddings of the query, the positive candidate, and any sample from the candidate set, respectively. is the temperature parameter. means the cosine similarity; More details are given in Appendix C. Throughout the ablation studies, we report the Average Recall as the evaluation metric, which is computed as the mean of Recall@5 or Recall@10 across the benchmark tasks (c.f. Appendix A). We then explore three major axes of design decisions in the following subsections:
@@ 3.1.1 Embedding Extraction
We first investigate the method for extracting embeddings of the token sequence of multimodal queries using MLLMs. Following prior approaches, the instruction is directly concatenated with the query. Existing methods can basically be divided into two categories based on the way embeddings are extracted. (1) Last token: previous works (Liu et al., 2024b; Jiang et al., 2024b; Zhang et al., 2024d; Jiang et al., 2024a; Gu et al., 2025) typically employ a prompt for compression. Generally, for multimodal input, the prompt is “<image><text>…<image><text>Summarize above image and sentence in one word: emb”, where <image> and <text> are the placeholders. In this way, the feature of last token (i.e., <emb>) is used as the output embedding of the input sequence. (2) Mean token: an alternative method, adopted by MM-Embed (Lin et al., 2024), transforms the unidirectional attention mechanism into a bidirectional one and applies mean pooling to the features from the last hidden states to derive the embedding for the entire sequence.
@@
| Bidirectional | Mean token | (“in one word”) | Avg. | Avg. | |
| 0 | Causal | Last token | 56.6 | 54.8 | |
| 1 | Bidirectional | Last token | 55.5 | 52.6 | |
| 2 | Causal | Mean token | 33.7 | 27.6 | |
| 3 | Bidirectional | Mean token | 51.7 | 46.1 | |
| 4 | Bidirectional | Mean token | 57.2 | 55.2 |
@@
In conclusion, generating embeddings using bidirectional attention combined with mean pooling from the sequence demonstrates superior performance compared to the commonly employed approach of combining compression prompts with the last token mechanism. This difference may stem from the last token embedding being affected by recency bias, leading to an excessive reliance on the output of the final token. Notably, this finding aligns with the conclusions of NV-Embed (Lee et al., 2024), yet diverges from those presented in GME (Zhang et al., 2024d), offering a distinct perspective on architectural design and performance optimization.
@@ 3.1.2 Instruction Integration
Motivated by prior studies on text embeddings (Lee et al., 2024; BehnamGhader et al., 2024), we mask out the instruction tokens during the mean pooling process, as these tokens have already influenced the output features through self-attention. Experimental results show that ID-0 achieves a 0.1%/0.3% improvement compared to ID-1 and 2.0%/12.1% improvement compared to ID-2 in Table 3. We hypothesize that this improvement arises primarily from that: the query inherently incorporates instruction information during forward propagation thought bidirectional self-attention mechanism. Therefore, by filtering out instruction tokens and focusing on feature comparisons between the query and candidate, the approach mitigates calculation bias. While the numerical improvement is modest, this step effectively eliminates theoretical instruction bias without compromising performance.
@@ 3.1.3 Progressive Transition
We set a baseline by vanilla instruction tuning on multimodal retrieval data from the M-BEIR dataset, presented in Table 3 under ID-0. Following prevalent approaches (Liu et al., 2024b; Jiang et al., 2024a), we use text-only retrieval data for pre-training, which improves performance on various retrieval tasks. Training on the NLI dataset (Gao et al., 2021) significantly enhances our model’s performance, as indicated by ID-1 in Table 3. Despite these improvements, a notable gap persists when transitioning from single-modality, single-instruction tasks to multimodal, multi-instruction tasks. To address this gap, we propose a progressive approach incorporating an additional pre-training step using text-image retrieval data from the CC3M dataset (Sharma et al., 2018). This step further enhances performance, as demonstrated by ID-2 in Table 3. Since LLM was originally trained with causal attention, switching to bidirectional attention degrades text-visual encoder alignment. We speculate that text-based fine-tuning strengthens the text encoder’s semantic representation, while image-text pairs further align the text-visual encoders. During experiments, we train the model with NLI using unidirectional InfoNCE, followed by CC3M using bidirectional InfoNCE. Furthermore, we further analyze the impact of training data selection in Appendix D, which reveals that the concise text in CC3M aligns better with retrieval tasks compared to other datasets (e.g., ShareGPT4V (Hurst et al., 2024) and TART (Asai et al., 2023)).
Therefore, we formally propose the progressive transition method, adopting a stepwise training strategy to adapt decoder-only MLLMs for embedding tasks. Following the principle of advancing from simpler to more complex tasks (Bengio et al., 2009), this phased approach ensures a smooth transition through three key steps: (1) Text Retrieval Adaptation: Learning retrieval tasks using text-only datasets to establish foundational capabilities. (2) Cross-modal Alignment: Advancing to cross-modal retrieval tasks through text-image paired data. (3) Instruction-tuned Multimodal Retrieval: Mastering instruction-guided multimodal retrieval tasks with comprehensive training on multimodal datasets.
@@ 3.2.1 Interactions Among Large Batch Size, Learning Rate and Temperature
Our experimental results dispute the widely held belief that “increasing batch size directly improves performance”, indicating that this assumption requires closer examination. By analyzing the results presented in Table 4, we identify three key findings as follows. (1) Experiments with ID-0, ID-2, and ID-4 indicate that increasing the batch size does improve model performance; however, the performance gains tend to plateau as the batch size continues to grow, as observed in comparisons between ID-5 and ID-6. (2) Consistent with prior research (Goyal et al., 2017) that highlights the necessity of scaling the learning rate (lr) alongside batch size, our contrastive learning experiments reveal similar trends. Specifically, simply increasing the batch size without adjusting the learning rate results in marginal improvements (ID-0 vs. ID-1). In contrast, applying scaling rule (ID-2) significantly boosts performance to 58.3%, which aligns with the findings from previous works. (3) Our investigation into the effects of temperature parameters reveals that employing a learnable temperature significantly enhances performance. The learnable mechanism effectively optimizes the sharpness of the probability distribution during training, superior to fixed temperature settings used in prior works such as LamRA (Liu et al., 2024b) and MM-Embed (Lin et al., 2024). Empirical results (comparing ID-4/ID-6 with ID-7/ID-8) demonstrate that this dynamic adjustment yields consistent performance gains (e.g., in Local Avg), while also exhibiting stable and adaptive convergence across different batch sizes. More detailed ablation study on learnable temperature is provided in Appendix D.
@@ 3.2.2 Continual Training with Hard Negative Mining
The initial random negative sampling yields basic discriminative ability but struggles with challenging false positives: negatives that exhibit high semantic similarity yet are incorrect matches. Recent studies have employed hard negative mining (Lin et al., 2024; Lan et al., 2025) or hard negative reweighting in loss computation (Hou and Li, 2023; Radenovic et al., 2023; Zhuang et al., 2024) to address this issue. In this work, we explore related settings. However, as shown in Table 5 for ID-0, directly selecting the top- hard negatives proved to be excessively challenging, resulting in complete failure. Simply setting to a larger value (e.g., 50) prevents model collapse but fails to yield any performance improvement. Additionally, incorporating in-batch negatives results in a performance decline compared to the baseline in this table, as observed by ID-1. We hypothesize that some hard negatives are false negatives, which can adversely affect model training. Specifically, naively maximizing the loss on all hard negatives often leads to model collapse because the model is forced to push away semantically valid candidates that are labeled as negatives due to dataset noise. To address this issue, we propose a straightforward filtering method that eliminates hard negatives with scores exceeding a predefined threshold. This approach led to a significant improvement, as demonstrated by ID-2.
We apply a hard negative mining strategy as follows: (1) Feature Extraction: Using the model from the previous training stage, we extract features for all queries and candidates, compute similarity scores (excluding positive matches), and rank candidates to identify the most challenging negatives per query; (2) Filtered Hard Negative: For each query, we filter out negative samples that exceed a predefined threshold (considered as false negatives) and then select the top- negative samples as hard negatives; (3) Balanced Training: To prevent model convergence difficulties from excessive hard negatives and to accelerate training, we select a suitable -value and mix these hard negatives with other in-batch negatives. We then perform continual finetuning by InfoNCE. In our experiments, we set and the threshold to 0.7, filtering out any negative samples with scores exceeding this limit. The choice of for hard negative mining strikes a balance between training efficiency and sample quality.